OpenStudy (anonymous):

Help me with this. The angle between \[3i-j\] and \[2i+\lambda j\] is \[\frac{pi}{4}\]. Find the value of \[\lambda\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

u figured it out ? :)

OpenStudy (anonymous):

\[(3i-j).(2i+\lambda j)=|3i-j|.|2i+\lambda j|\space cos\frac{\pi}{4} \] \[6-\lambda =\sqrt{10}.\sqrt{4+\lambda^2}\times \frac{1}{\sqrt{2}}\] \[6-\lambda =\sqrt{5}.\sqrt{4+\lambda^2}\] \[(6-\lambda)^2 =5.(4+\lambda^2)^2\]

OpenStudy (anonymous):

Am I correct ?

ganeshie8 (ganeshie8):

yes looks good, solve it all the way

OpenStudy (anonymous):

ok :)

ganeshie8 (ganeshie8):

last step, i see a mistake it should be :- \((6-\lambda)^2 =5.(4+\lambda^2)\)

OpenStudy (anonymous):

Ohh Yeah :)

OpenStudy (anonymous):

Thanks for showing me.

ganeshie8 (ganeshie8):

np :) since its a quadratic, u wil get two values for lambda

OpenStudy (anonymous):

\[\lambda^2-12\lambda+36=20+5\lambda^2\] \[4\lambda^2+12\lambda-16=0\] \[\lambda^2+3\lambda-4=0\] \[\lambda^2+4\lambda-\lambda-4=0\] \[(\lambda+4)(\lambda-1)=0\] \[\lambda=-4 \space or \space \lambda=1\]

OpenStudy (anonymous):

Hurray I got it :)

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

Just bothered you with a mistake of mine.

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):